Please help me solve:
find the solution to the equation graphically.
1. 5x - 3 = x - 5
2. 3x + 5 = 4(x - 4) + 4
Thank you.

Answer by checkley77(12844)

1. 5x - 3 = x - 5
5x-x=-5+3
4x=-2
x=-2/4
x=-1/2 ans.
Proof:
5(-1/2)-3=1/2-5
-5/2-3=(1-5*2)/2
(-5-3*2)/2=(1-10)/2
(-5-6)/2=-11/2
-11/2=-11/2
-------------------------------
2. 3x + 5 = 4(x - 4) + 4
3x+5=4x-16+4
3x-4x=-16+4-5
-x=-17
x=17 ans.
Proof:
3*17+5=4(17-4)+4
51+5=4*13+4
56=52+4
56=56